On dividing 7 by 2/5, the result is (a) 14/2 (b) 35/4 (c) 14/5 (d) 35/2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (b) is correct here
= 7/ (2/5)
= 7 × (5/2)
= 35/2


Thanks